Carole is an active 55-year old, who is classified as being in an average fitness category,...

  1. Carole is an active 55-year old, who is classified as being in an average fitness category, with a resting HR of 80 bpm.  Calculate her target heart rate range for an intensity of 60-85% of predicted max heart rate using the Karvonen heart rate reserve method.  (2 points)

  1. Marvin wants to lose weight and is exercising at a pedal rate of 60 rpms at 60% of his VO2max on a Monark Cycle. What would the resistance on his cycle be? His Max VO2 is 42 ml/kg/min and he weighs 200 lb.

(3 points)
Answer #1

Answer:

Based on the given information:

Part I:

  • Carole - Resting HR = 80 bpm
  • Carole age is 55 years
  • Target heart rate range for an intensity of 60-85% of predicted max heart rate
  • The formulae used as per Karvonen heart rate reserve method is:
  • Target Heart Rate = [(max HR − resting HR) × %Intensity] + resting HR
  • Maximum heart rate can be determined using a simple formula: 220 - age = 220 - 55 = 165
  • For intensity 60%, Target Heart Rate = [(165 −80) × 60%] + 80 = 131 bpm
  • For intensity 85%, Target Heart Rate = [(165 −80) × 85%] + 80 = 152 bpm
  • Therefore, Low HR is 131 bpm and High HR is 152 bpm

Part II:

  • Pedal rate of 60 rpms at 60% of his VO2max on a Monark Cycle.
  • Max VO2 is 42 ml/kg/min and
  • Marvin Weight = 200 lb = 90.7 Kg
  • Resistance on his cycle be?
  • One revolution of the flywheel is = distance of 6.0 meters
  • VO2 Cycling = (1.8 x WR)/BW + 3.5 + 3.5  
  • 42 = (1.8*WR)/90.7 + 3.5 + 3.5
  • WR = (42 - 3.5 - 3.5)*90.7 / 1.8 = 1763.6 kgm
  • Monark Cycle Work Rate: Resistance X Revs/min x 6m/rev = Resistance * 60 * 6 = Resistance *360
  • WR = 1763.6 = Resistance *360
  • Therefore, Resistance = 1763.6 / 360 = 4.9 Kg